How would u find a line that is parallel to y=2/3x+5, passing through (4,-5)
Found 2 solutions by Marieanne, Alan3354:
Answer by Marieanne(1)

I will hopefully help you I will try my best. So it says find the line that is parallel to y=2/3x+5. So the slope in this situation is 2/3 so rise over run and says its parallel the slope stays the same as it is parallel so then just the y value changes(b in y=mx+b form) then since the y value they ask for is -5 as it passes through that point plot that dot on a graph first and continue the same slope in both cases slope same y-int different.
I am just a student as well I just attempted to help another fellow student out.

Answer by Alan3354(69443)

How would u find a line that is parallel to y=2/3x+5, passing through (4,-5)
----------
Since the lines are parallel, the slopes are equal, = 2/3
------
y+5 = (2/3)*(x-4)
3y + 15 = 2x - 8
2x - 3y = 23